# Real-time Surgical Instrument Segmentation in Video Using Point Tracking and Segment Anything

### v1

@@ Abstract
The Segment Anything Model (SAM) is a powerful vision foundation model that is revolutionizing the traditional paradigm of segmentation. Despite this, a reliance on prompting each frame and large computational cost limit its usage in robotically assisted surgery. Applications, such as augmented reality guidance, require little user intervention along with efficient inference to be usable clinically. In this study, we address these limitations by adopting lightweight SAM variants to meet the speed requirement and employing fine-tuning techniques to enhance their generalization in surgical scenes. Recent advancements in Tracking Any Point (TAP) have shown promising results in both accuracy and efficiency, particularly when points are occluded or leave the field of view. Inspired by this progress, we present a novel framework that combines an online point tracker with a lightweight SAM model that is fine-tuned for surgical instrument segmentation. Sparse points within the region of interest are tracked and used to prompt SAM throughout the video sequence, providing temporal consistency. The quantitative results surpass the state-of-the-art semi-supervised video object segmentation method on the EndoVis 2015 dataset, with an over 25 FPS inference speed running on a single GeForce RTX 4060 GPU.
Keywords:
Tracking Any Point Segment Anything Surgical Instrument Segmentation.
@@ 1 Introduction
Despite SAM’s strong automatic mask generation ability, achieving expected segmentation results in practice often requires appropriate prompts. Providing specific points or descriptive text of the target object can significantly improve the segmentation accuracy. Leveraging the long-term tracking capabilities of the Tracking Any Point (TAP) models, we employ an online point tracker, CoTracker [9], to provide sparse point prompts for SAM. Similar to SAM-PT [8] and DEVA [23], our pipeline decouples video object segmentation (VOS) into image-level segmentation, which can be task-specific, and a universal temporal propagation. Compared to end-to-end VOS, our “tracking-by-detection” framework can take full advantage of smaller image-level datasets via fine-tuning a task-specific image segmentation model and using it in tandem with a point tracker to maintain temporal consistency.
In summary, our contribution is threefold: (1) we present a real-time video surgical instrument segmentation framework that achieves superior segmentation performance and is suitable for clinical usage due to its good efficiency; (2) we investigate the point prompt-based fine-tuning strategy (will open source) for lightweight SAM using surgical datasets, and the model fine-tuned on only two datasets shows promising generalization on unseen surgical videos; (3) a naive combination of SAM and TAP cannot achieve ideal performance from both the segmentation accuracy and the inference efficiency perspectives for online SIS, while our presented pipeline significantly enhances the performance on both fronts.
@@ 2 Related Work
2.0.1 Tracking Any Point:
The goal of Tracking Any Point (TAP) is to estimate the motion of arbitrary physical points throughout a video. TAP-Vid [12] first formalized this task alongside a benchmark dataset and baseline method for TAP. Recent work has showcased the promising online TAP capability and exhibited great robustness to occlusion and exit from the field of view. PIPs++ [13] and TAPIR [14] demonstrate a substantial robustness under occlusion and achieve real-time inference speed on high-resolution video. Notably, CoTracker [9] achieves state-of-the-art tracking performance by jointly tracking a set of query points. CoTracker is an online algorithm that processes video sequentially through a sliding window. Optical flow [10, 11] can be used for TAP but tends to accumulate errors over time and faces challenges in handling occlusions, which are common occurrences in surgical scenarios.
2.0.2 Segment Anything:
@@ 3.1 Pre-processing
We use the SAM model to generate the initial mask by simply inputting a few points or bounding boxes. Inspired by recent advancements in vision-language models, several zero-shot segmentation models based on text prompts have been developed [24, 25]. To achieve the “fully” automatic pipeline, we incorporate a text-promptable segmentation model CLIPSeg [24] to automatically generate the initial mask by setting a text prompt, “surgical tool”. While CLIPSeg can only provide a coarse initial mask, it remains feasible for query point selection as long as its output roughly covers the region of the target instrument.
We initialize query points using the first frame mask. We investigate various query point sampling strategies, including random sampling, uniform sampling on a grid, SIFT keypoints, Shi-Tomasi corner points, and K-Medoids [35] clustering centers. Further details on these sampling strategies are provided in the Supplementary Material. We selected K-Medoids clustering centers because they ensure even partitioning of the entire cluster. The number of medoids assigned to each instance ranges from 1 to 9, and we choose 5 in the experiments. Let
@@ 3.3 Fine-tuning the Segment Anything Model
As for the segmentation model, we first tested two state-of-the-art lightweight SAM variants, MobileSAM and Light HQ-SAM [6]. However, we observed that both methods perform poorly in situations where specularity, blood, or weak lighting is present, as shown in Fig. 4. Driven by this limitation, enhancing the generalization of the lightweight SAM for surgical instrument segmentation becomes crucial. The state-of-the-art MedSAM [15] has demonstrated that fully fine-tuning SAM for medical images can yield promising results. Fully fine-tuning refers to freezing the prompt encoder and updating both the image encoder and mask decoder. Compared with strategies that only update the mask decoder or introducing the adapter layer, fully fine-tuning achieves superior performance.
@@
We used the AdamW [32] optimizer for training, with a batch size of 32. The initial learning rate is set to 1e-5 and follows a cosine decay schedule. All images are resized to 1024 by 1024 and undergo random up-down and left-right flip data augmentation.
4 Experiments and Results
4.0.1 Dataset:
We conduct quantitative comparisons on the EndoVis 2015 [30] and the UCL dVRK [28] datasets to validate the feasibility and performance of our proposed framework. Additionally, we display qualitative results on these two datasets as well as results on the unlabeled high-resolution STIR [29] dataset. The EndoVis 2015 dataset provides annotations for both rigid and articulated robotic instruments in laparoscopic surgery. In this study, we focus solely on the robotic instruments. For both datasets, we keep the original data split for fine-tuning and testing. The frame rates of the videos and annotations in the EndoVis 2015 is 25 FPS. In the UCL dVRK dataset, images are collected and annotated at 6.67 FPS. In the STIR dataset, the video frame rate is 30 FPS.
4.0.2 Quantitative Results:
We perform extensive comparisons between different SAM variants and the state-of-the-art semi-supervised VOS method, XMem [31]. We adopt the standard VOS metric [33], $\mathcal{J}\&\mathcal{F}$, the average of Jaccard region similarity ($\mathcal{J}$) and contour accuracy ($\mathcal{F}$) for quantitative comparison. Table 1 and 2 show the quantitative results of the EndoVis 2015 and UCL dVRK datasets, respectively. $\mathcal{J}$ and $\mathcal{F}$ results are provided in the Supplementary Material. The inference speed of our fine-tuned MobileSAM is about 40 milliseconds, while the ViT-H SAM’s inference speed is about 0.9 seconds. The CoTracker’s frame rate fluctuates in the 50-60 FPS range. All experiments are running on a single GeForce RTX 4060 GPU.
Table 1: Quantitative Results ($\mathcal{J}\&\mathcal{F}$) on the EndoVis 2015 Dataset. The abbreviations PT and FF-SAM stand for point tracker and fully fine-tuned MobileSAM, respectively. The same abbreviations are used in Table 2.
Method Video 1 Video 2 Video 3 Video 4 Video 5 Video 6
PT+MobileSAM [7] 79.9 87.8 84.7 59.5 65.3 74.9
PT+Light HQ-SAM [6] 81.3 88.7 80.1 58.8 63.5 75.1
PT+ViT-H SAM [4] 87.2 89.3 86.9 81.2 77.0 81.9
PT+FF-SAM 89.0 92.1 88.9 84.6 86.3 85.5
XMem [31] 87.3 91.7 88.7 84.2 83.1 83.8
Table 2: Quantitative Results ($\mathcal{J}\&\mathcal{F}$) on the UCL dVRK Dataset.
Method Video 11 Video 12 Video 13 Video 14
PT+MobileSAM [7] 55.6 33.6 50.6 64.3
PT+Light HQ-SAM [6] 36.6 70.6 60.7 80.2
PT+ViT-H SAM [4] 73.6 81.2 79.6 80.8
PT+FF-SAM 89.9 89.5 94.0 91.1
XMem [31] 93.5 93.4 94.4 93.1
Figure 4: Visualization of qualitative results from our proposed methods on different datasets. The point tracker is CoTracker, which we pair with different SAM variants. The red dots are the query points for prompting SAM. The images are from STIR, EndoVis 2015, and UCL dVRK, respectively (top to bottom).
4.0.3 Qualitative Results:
@@
Fig. 4 displays a few qualitative results of various methods on different datasets. To demonstrate the generalization of the proposed method, we test our framework, which is fine-tuned using the EndoVis 15 and UCL dVRK dataset, on some videos from the STIR dataset. Note that the model has never seen this dataset.
@@ 5 Discussion
Overall, our proposed method outperforms the state-of-the-art semi-supervised VOS model, XMem, on the EndoVis 2015 dataset. Notably, in the one-minute long videos, Video 5 and 6, our proposed method exhibits obvious improvement compared to XMem. This shows that our framework can robustly leverage the temporal information provided by universal TAP methods. The significance of fine-tuning is demonstrated by the substantial improvement compared to non-fine-tuned SAMs, even that of the powerful ViT-H SAM. Note that the inference speed of the ViT-H SAM is far away from real-time. In the UCL dVRK dataset, our proposed method also achieves promising results, although it does not outperform XMem. All four UCL dVRK testing videos are captured under very weak illumination conditions, which is unrealistic in real surgery. The dark scenes make it challenging for SAM to distinguish the boundary from the tissue background, especially when there are no such weak lighting conditions in the training set. In contrast, the mask propagation-based XMem takes the full first frame ground truth for its inference. The low video frame rate (6.67 FPS) of the UCL dVRK dataset also poses an obstacle for the TAP, thereby making it hard to provide effective point prompts throughout the video sequence. Compared to XMem, our method has other advantages. XMem requires an accurate first frame mask, while our pipeline only needs a known text prompt “surgical tool”. Furthermore, XMem cannot recognize new objects during the video, while our method can tackle this by enabling users to easily pick a few new query points, instead of providing a high-quality mask. The quantitative results and qualitative examples depicted in Fig. 4 show the superiority of the fine-tuned MobileSAM over other lightweight SAM variants and even the ViT-H SAM.
6 Conclusion
@@
In this study, we present a novel framework using a universal TAP and a fine-tuned lightweight SAM for real-time surgical instrument segmentation in video. Its commendable efficiency and accuracy makes it suitable for applications in clinical settings. We investigate the availability of fine-tuning MobileSAM using point prompts and demonstrate the importance of fine-tuning for SIS. Extensive experiments validate the advancement of our proposed pipeline. Furthermore, our SAM + TAP pipeline demonstrates the potential to serve as a strong VOS baseline by integrating other image segmentation models.

### v2

# Augmenting Efficient Real-time Surgical Instrument Segmentation in Video with Point Tracking and Segment Anything

@@ Abstract
The Segment Anything Model (SAM) is a powerful vision foundation model that is revolutionizing the traditional paradigm of segmentation. Despite this, a reliance on prompting each frame and large computational cost limit its usage in robotically assisted surgery. Applications, such as augmented reality guidance, require little user intervention along with efficient inference to be usable clinically. In this study, we address these limitations by adopting lightweight SAM variants to meet the efficiency requirement and employing fine-tuning techniques to enhance their generalization in surgical scenes. Recent advancements in Tracking Any Point (TAP) have shown promising results in both accuracy and efficiency, particularly when points are occluded or leave the field of view. Inspired by this progress, we present a novel framework that combines an online point tracker with a lightweight SAM model that is fine-tuned for surgical instrument segmentation. Sparse points within the region of interest are tracked and used to prompt SAM throughout the video sequence, providing temporal consistency. The quantitative results surpass the state-of-the-art semi-supervised video object segmentation method XMem on the EndoVis 2015 dataset with 84.8 IoU and 91.0 Dice. Our method achieves promising performance that is comparable to XMem and transformer-based fully supervised segmentation methods on ex vivo UCL dVRK and in vivo CholecSeg8k datasets. In addition, the proposed method shows promising zero-shot generalization ability on the label-free STIR dataset. In terms of efficiency, we tested our method on a single GeForce RTX 4060/4090 GPU respectively, achieving an over 25/90 FPS inference speed. Code is available at: https://github.com/wuzijian1997/SIS-PT-SAM
@@ 1 Introduction
Despite SAM’s strong automatic mask generation ability, achieving expected segmentation results in practice often requires appropriate prompts. Providing specific points or descriptive text of the target object can significantly improve the segmentation accuracy. Leveraging the long-term tracking capabilities of the Tracking Any Point (TAP) models, we employ an online point tracker, CoTracker [9], to provide sparse point prompts for SAM. Similar to SAM-PT [8] and DEVA [24], our pipeline decouples video object segmentation (VOS) into image-level segmentation, which can be task-specific, and a universal temporal propagation. Compared to end-to-end VOS, our “tracking-by-detection" framework can take full advantage of smaller image-level datasets via fine-tuning a task-specific image segmentation model and using it in tandem with a point tracker to maintain temporal consistency.
In summary, our contribution is twofold: (1) we present a real-time video surgical instrument segmentation framework that achieves superior segmentation performance and is suitable for clinical usage due to its good efficiency; (2) we investigate the point prompt-based fine-tuning strategy (will open source) for lightweight SAM using surgical datasets, and the model fine-tuned on only two datasets shows promising generalization on unseen surgical videos.
@@ 2 Related work
The goal of Tracking Any Point (TAP) is to estimate the motion of arbitrary physical points throughout a video. TAP-Vid [12] first formalized this task alongside a benchmark dataset and baseline method for TAP. Recent work has showcased the promising online TAP capability and exhibited great robustness to occlusion and exit from the field of view. PIPs++ [13] and TAPIR [14] demonstrate substantial robustness under occlusion and achieve real-time inference speed on high-resolution video. Notably, CoTracker [9] achieves state-of-the-art tracking performance by jointly tracking a set of query points. CoTracker is an online algorithm that processes video sequentially through a sliding window. Optical flow [10, 11] can be used for TAP but tends to accumulate errors over time and faces challenges in handling occlusions, which are common occurrences in surgical scenarios.
@@ 3.1 Pre-processing
We use the SAM model to generate the initial mask by simply inputting a few points or bounding boxes. Inspired by recent advancements in vision-language models, several zero-shot segmentation models based on text prompts have been developed [25, 26]. To achieve the “fully" automatic pipeline, we incorporate a text-promptable segmentation model CLIPSeg [25] to automatically generate the initial mask by setting a text prompt, “surgical tool". While CLIPSeg can only provide a coarse initial mask, it remains feasible for query point selection as long as its output roughly covers the region of the target instrument.
We initialize query points using the first frame mask. We investigate various query point sampling strategies, including random sampling, uniform sampling on a grid, SIFT keypoints, Shi-Tomasi corner points, and K-Medoids [36] clustering centers. We selected K-Medoids clustering centers because they ensure even partitioning of the entire cluster. The number of medoids assigned to each instance ranges from 1 to 9, and we choose 5 in the experiments. Let
@@ 3.3 Fine-tuning the Segment Anything Model
As for the segmentation model, we first tested two state-of-the-art lightweight SAM variants, MobileSAM and Light HQ-SAM [6]. However, we observed that both methods perform poorly in situations where specularity, blood, or weak lighting is present, as shown in Fig. 3. Driven by this limitation, enhancing the generalization of the lightweight SAM for surgical instrument segmentation becomes crucial. The state-of-the-art MedSAM [15] has demonstrated that fully fine-tuning SAM for medical images can yield promising results. Fully fine-tuning refers to freezing the prompt encoder and updating both the image encoder and mask decoder. Compared with strategies that only update the mask decoder or introduce an adapter layer, fully fine-tuning achieves superior performance.
@@
We used the AdamW [33] optimizer for training, with a batch size of 32. The initial learning rate is set to 1e-5 and follows a cosine decay schedule. All images are resized to 1024 by 1024 and undergo random up-down and left-right flip data augmentation. All images are Min-Max normalized and standardized.
4 Results
4.1 Datasets
We conduct quantitative comparisons on the EndoVis 2015 [31], UCL dVRK [29] and CholecSeg8k [38] datasets to validate the feasibility and performance of our proposed framework. We finetune the SAM model using the training set of the ROBUST-MIS 2019 [39] dataset and provide qualitative results because there is no video-level annotation for quantitative evaluation. To test the zero-shot generalization, we test our method on the unlabeled STIR [30] dataset and display qualitative results.
The EndoVis 2015 dataset provides 25 FPS videos and corresponding articulated da Vinci robotic instruments in ex vivo background. It consists of four 45-second training videos and six testing videos (four 15-second and two 60-second videos). Note that there is only one type of instrument (needle driver) in the training set while two types (needle driver and scissor) in the testing set. We keep the original data split for fine-tuning and testing.
The ex vivo UCL dVRK dataset consists of 14 videos of 300 frames with corresponding binary segmentation masks. The dataset is split into training (Video 1-8), validation (Video 9 and 10) and testing (Video 11-14) sets. The videos are collected and annotated at 6.7 FPS.
Based on the Cholec80 dataset [40], the CholecSeg8k dataset consists of 8080 frames (from 17 in vivo cholecystectomy video clips) with segmentation ground truth. We choose 8 consecutive clips, which are 905 frames in total, as the testing set. The rest of the data is split into the training and validation set as 80% and 20% ratio.
We select two extra datasets for qualitative evaluation. The ROBUST-MIS 2019 data comprises 10,040 annotated images, of which 5,983 are in the training set. The testing set is divided into three stages with increasing levels of difficulty. The dataset provides one labeled frame per clip. The STIR dataset consists of high-resolution label-free videos collected by da Vinci Xi.
4.2 Quantitative Results
We perform comparisons between the state-of-the-art semi-supervised VOS method XMem [32], and fully supervised image segmentation methods TransUNet [22] and SwinUNet [37]. The ablation study of different SAM variants (MobileSAM, HQ-SAM Light, and the default SAM) demonstrates the significant performance improvement of finetuning. All these SAMs use the CoTracker as the online point tracker. We adopt the widely-used segmentation metrics, $IoU$ and $Dice$, for quantitative comparison and ablation study. Table 1 displays the quantitative results on three datasets, in which the first 3 rows are comparison results, the next 3 rows are the ablation results, and the bottom row is ours.
Table 1: Quantitative Results of Different Datasets
Methods EndoVis 2015 [31] UCL dVRK [13] CholecSeg8k [38]
IoU Dice IoU Dice IoU Dice
TransUNet [22] 57.7 71.8 79.1 88.0 71.7 82.8
SwinUNet [37] 59.5 73.0 81.1 89.3 81.5 89.5
XMem [32] 82.6 89.3 91.9 95.4 81.6 87.9
PT+MobileSAM [7] 70.1 80.6 45.0 57.1 49.7 57.4
PT+HQ-SAM Light [6] 69.0 80.2 55.9 68.1 61.3 70.4
PT+SAM (ViT-H) [4] 79.6 88.3 74.3 83.3 64.5 70.6
Ours 84.4 91.0 89.4 93.8 81.9 88.6
Figure 4: Visualization of segmentation results from several methods on the EndoVis 2015 dataset, in which the images are acquired from the testing Video 1, 5, and 6, respectively (from top row to bottom row). Note that red dots in Fig. 4 - 8 are the point prompts tracked by CoTracker.
Figure 5: Visualization of segmentation results from several methods on the UCL dVRK dataset, in which the images are acquired from the testing Video 1, 3, and 4, respectively (from top row to bottom row).
Figure 6: Segmentation results of different methods on the CholecSeg8k dataset. The y-axis is along the frame order (10th, 20th, 50th, and 70th) in one clip.
Figure 7: Qualitative results from our proposed methods on the ROBUST-MIS 2019 dataset, in which columns from left to right are the samples from stages 1, 2, and 3, respectively.
4.3 Qualitative Results
@@
Fig. 4 and Fig. 5 display a few predicted masks from various methods on EndoVis 2015 and UCL dVRK datasets. Fig. 6 is the segmentation results of the frame 10, 20, 50, and 70, which are sampled on a video clip from the testing set of the CholecSeg8k dataset. Fig. 7 is the visualization of segmentation results on different difficulty testing stages of the ROBUST-MIS 2019 dataset. To evaluate the generalization of our method, we test our framework, which is fine-tuned using the EndoVis 15 and UCL dVRK dataset, on some videos from the STIR dataset. Fig. 8 shows some segmentation results of our method, in which frame 397 is the end frame. The segmentation performance of our method is robust in this video.
Figure 8: Visualization of qualitative results from our proposed methods on STIR dataset.
4.4 Efficiency
Table 2 displays the metrics of efficiency of different methods. The inference latency, inference memory, and learnable parameters represent the time efficiency, computation efficiency for inference, and computation efficiency for training. We run the inference procedure of each method on a laptop with RTX 4060 GPU and a desktop with RTX 4090 GPU. The inference latency of our method achieves 38 ms (26 FPS) and 11 ms (90 FPS), which are real-time and over real-time inference speed, on these two machines. The CoTracker’s inference speed fluctuates in the 50-60 FPS range. In terms of computational efficiency, our proposed method only requires 2.8G inference memory. Our method is also training efficient with a small 10.1M parameters. We only list learnable parameters of partial methods, because the other methods do not involve training in this study.
Table 2: Efficiency Metrics with RTX 4090 and 4060 GPUs
Methods Inference Inference Learnable
Latency (ms) Memory (G) Param. (M)
4090 4060
XMem [32] 4 5 2.0 -
TransUNet [22] 7 8 2.8 105.3
SwinUNet [37] 4 7 1.9 27.2
PT+MobileSAM [7] 11 38 2.8 -
PT+Light HQ-SAM [6] 12 40 2.8 -
PT+ViT-H SAM [4] 220 1300 8.2 -
Ours 11 38 2.8 10.1
@@ 5 Discussion
Overall, our method outperforms the state-of-the-art semi-supervised VOS model, XMem, on the EndoVis 2015 dataset. Notably, in the one-minute videos, Video 5 and 6, our proposed method exhibits obvious improvement compared to XMem. This shows that our framework can robustly leverage the temporal information provided by universal TAP methods. The TransUNet and SwinUNet cannot recognize the scissors in testing videos because of the overfitting to the instrument type in the training videos.
@@
As for the UCL dVRK dataset, our proposed method achieves promising performance with slightly weaker quantitative results than XMem. All four UCL dVRK testing videos are captured under weak illumination conditions, which is unrealistic in real surgery. The dark scenes make it challenging for SAM to distinguish the boundary from the tissue background, especially when there are no such weak lighting conditions in the training set. In contrast, the mask propagation-based XMem takes the full first frame ground truth for. The low video frame rate (6.7 FPS) of the UCL dVRK dataset also poses an obstacle for the TAP, thereby making it hard to provide effective point prompts throughout the video sequence.
The quantitative results on the in vivo CholecSeg8k dataset achieve the best $IoU$ and second best $Dice$. The segmentation performance of our method is comparable to the state-of-the-art fully supervised image segmentation model SwinUNet.
The ablation study demonstrates the significance of fine-tuning by the substantial improvement compared to non-fine-tuned SAMs, even that of the powerful SAM with ViT-H backbone. Note that the inference speed of the ViT-H SAM is far away from real-time.
Compared to the propagation-based model such as XMem, our method has other advantages. XMem requires an accurate first frame ground truth for inference, while our pipeline only needs a known text prompt “surgical tool". Furthermore, XMem cannot recognize new objects during the video, while our method can tackle this by enabling users to easily pick a few new query points, instead of providing a high-quality mask.
6 Limitations and Future Work
In general, our method achieves satisfied segmentation performance along with good efficiency in extensive SIS video datasets. However, when processing more challenging datasets like SAR-RARP [41], the segmentation performance of our method is not satisfactory, as depicted in Fig. 9. The videos of the SAR-RARP dataset are recorded during real robot-assisted laparoscopic prostatectomy with a cluttered scene, significant blood, specular reflection, wide-range rapid instrument motion, and frequent camera focus changes. These natures cause huge challenges for both point tracking and segment anything model.
To enhance the performance in more challenging surgical scenarios, we plan to design a novel SAM prompt method to implicitly leverage the spatio-temporal information from point tracking instead of the naive combination. Kinematics data is a strong prior for instrument identification. However, this is only available for robotic surgical instruments. Other modality information such as text description of surgical instruments can be leveraged as a prior knowledge.
Figure 9: Failure cases on the SAR-RARP dataset.
7 Conclusions
@@
In this study, we present a novel framework using a universal TAP and a fine-tuned lightweight SAM for real-time surgical instrument segmentation in video. Its commendable efficiency and accuracy make it suitable for applications in clinical settings. We investigate the availability of fine-tuning MobileSAM using point prompts and demonstrate the importance of fine-tuning for SIS. Extensive experiments validate the advancement of our proposed pipeline. Furthermore, our SAM + TAP pipeline demonstrates the potential to serve as a strong VOS baseline by integrating other image segmentation models.